OpenStudy (anonymous):

A plane landing on a small tropical island has just 63 m of runway on which to stop.If its initial speed is 61 m/s, what is the maximum acceleration of the plane during landing, assuming it to be constant? and please explain

ganeshie8 (ganeshie8):

you can use the equation : \(v^2-u^2 = 2as\)

OpenStudy (anonymous):

thanks really helped

ganeshie8 (ganeshie8):

yw !! did you get a = 61^2 / 2*63 ?

OpenStudy (anonymous):

yeah i did and another question how long does it take for the plane to stop with that acceleration? do i use a diffrent formula or ...

OpenStudy (dumbcow):

hmm don't forget "a" should be negative since plane is decelerating for time, use velocity equation v(t) = at + 61 ....set v=0 and solve for t

OpenStudy (anonymous):

ok thank i got -2.065573770488029

ganeshie8 (ganeshie8):

thats right, no -ve sign for time... its +2.0655

OpenStudy (anonymous):

yeah thanks

ganeshie8 (ganeshie8):

np.. :D